A research scientist wants to know how many times per hour a certain strand of bacteria reproduces. The mean is found to be 10.2 reproductions and the population standard deviation is known to be 2.1. If a sample of 1731 was used for the study, construct the 99% confidence interval for the true mean number of reproductions per hour for the bacteria. Round your answers to one decimal place.
Answer #1
sample mean 'x̄= 10.2
sample size    n= 1731.00
std deviation σ= 2.100
std errror ='σx=σ/√n= 0.0505
for 99 % CI value of z= 2.576
margin of error E=z*std error = 0.13
lower bound=sample mean-E= 10.1
Upper bound=sample mean+E= 10.3
from above 99% confidence interval for population mean =(10.1,10.3)